# Video LLMs for Temporal Reasoning in Long Videos

### v3

@@ 5.4 Qualitative Results
In this section, we discuss some qualitative results.
5.4.1 Dense Video Captioning Results
Fig. 4 shows example results of dense video captioning in the zero-shot setting on a YouCook2 [58] video. TimeChat [35] fails to predict any correct timestamps and tends to hallucinate. LongVLM [49] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions. In contrast, TemporalVLM produces considerably more accurate timestamps and captions. Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long video understanding.
5.4.2 Temporal Action Segmentation Results
Fig. 5 illustrates example results of temporal action segmentation in the supervised setting on an IndustryASM video. TimeChat [35] falls short in segmenting the video correctly and hallucinates actions not present in the video. LongVLM [49] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video. In contrast, TemporalVLM predicts action segments that are significantly closer to the ground truth.
@@ 6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. The extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes. Our dataset is recorded on factory floors and includes action and timestamp labels provided by industrial engineers for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on TimeIT [35] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
@@ Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.4 respectively. Next, Sec. A.5 compares the sizes of our model, TimeChat [35], and LongVLM [49], while Sec. A.6 evaluates our model for general video understanding. We then present several qualitative results in Sec. A.7, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Sec. A.8.
@@ A.4 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. The output of the clip encoder contains a time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
@@
Q-Former window size 32
Stride 32
Nnumber of visual tokens per window 32
@@
A.5 Comparison on Model Sizes
@@
TemporalVLM (Ours) 254 millions
@@
A.6 General Video Understanding Results

### v4

@@ 5.4 Qualitative Results
Dense Video Captioning Results. Fig. 4 shows example results of dense video captioning in the zero-shot setting on a YouCook2 [59] video. TimeChat [36] fails to predict any correct timestamps and tends to hallucinate. LongVLM [50] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions. In contrast, TemporalVLM produces much more accurate timestamps and captions. Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long videos.
Temporal Action Segmentation Results. Fig. 5 illustrates example results of temporal action segmentation in the supervised setting on an IndustryASM video. TimeChat [36] falls short in segmenting the video correctly and hallucinates actions not present in the video. LongVLM [50] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video. In contrast, TemporalVLM predicts action segments significantly closer to the ground truth.
@@ 6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts fused time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. The extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on TimeIT [36] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
@@ Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.4 respectively. Next, we provide a comparison with specialized models such as Vid2Seq [53] and QD-DETR [30] in Sec. A.5, followed by demonstrating the effects of number of training epochs in Sec. A.6. Sec. A.7 compares the sizes of our model, TimeChat [36], and LongVLM [50], while Sec. A.8 assesses the impact of number of clips. Sec. A.9 evaluates our model for general video understanding. We then present several qualitative results in Sec. A.10, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Sec. A.11.
@@ A.4 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q$ and overlap $o$. We then fuse the concatenated window-wise outputs of the Q-former via multiheaded self attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former. The output of the clip encoder contains a fused time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
@@
Q-Former window size $q$ 32
Overlap $o$ 16
Number of visual tokens per window 32
@@
A.5 Comparison With Specialized Models
Tab. 8 shows generalist models underperform specialized models such as Vid2Seq  [53] and QD-DETr [30] whose advantages come from task-specific training data and model designs. In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR  [30] leads to superior performance in highlight detection tasks. In contrast our TemporalVLM is trained on a simpler language modeling loss on lesser epochs. Generalist models such as ours exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
Model DVC VHD TVG
SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
Vid2Seq 7.9 47.1 27.3
QD-DETR 38.9 62.4
QD-DETR w/ Audio 39.0 62.9
MMN 50.5 29.7
VDI 52.3 31.4
TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 8: Comparisons with specialized models.
A.6 Effects of Number of Training Epochs
Tab. 9 shows the performance of our models across different training epochs.Since our model uses a learnable fusion layer in the clip encoder and a BiLSTM module to aggregate clip level features into a global representation, we observe the best results by training for 7 epochs. However, Tab. 9 reports our results with 3 epochs, which underperform those with 7 epochs but still outperform those of TimeChat and LongVLM.
Model DVC VHD TVG
SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
LongVLM (3 epochs) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat (3 epochs) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM (3 epochs) 3.2 12.9 19.7 23.9 42.3 50.0 25.9
TemporalVLM (7 epochs) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 9: Effects of number of training epochs.
A.7 Comparison on Model Sizes
@@ A.8 Impacts of Number of Short-Term Clips
TemporalVLM (Ours) 255 millions
@@
# Clips SODA_c CIDEr F1
2 2.6 9.2 18.5
4 3.1 11.3 18.9
6 3.4 13.2 20.0
A.9 General Video Understanding Results